# Variable Selection in Convex Piecewise Linear Regression

### v1

@@ 5 Discussion
$\leq 4\sum_{j=1}^{k}\Bigg{\{}\frac{\epsilon_{\mathrm{min}}}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}\bigg{[}\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}^{t}_{j}\|_{2}+\|\bm{\widetilde{\Pi}}_{\varphi^{t}\backslash\varphi^{t+1}}\bm{h}^{t}_{j}\|_{2}+4\lambda(k-1)\frac{\epsilon_{\mathrm{min}}}{\sqrt{k}}\|\bm{h}^{t}_{j}\|_{2}+\tau_{\mathrm{noise}}/\epsilon_{\mathrm{min}}\bigg{]}\Bigg{\}}^{2}$
$\leq 4\left[\frac{\epsilon_{\mathrm{min}}}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}\right]^{2}\sum_{j=1}^{k}\left\{\bigg{[}\sqrt{2}+4\lambda(k-1)\frac{\epsilon_{\mathrm{min}}}{\sqrt{k}}\bigg{]}\|\bm{h}^{t}_{j}\|_{2}+\tau_{\mathrm{noise}}/\epsilon_{\mathrm{min}}\right\}^{2}$
$\leq\underbrace{2\left\{\frac{\epsilon_{\mathrm{min}}}{(1-\varrho)\pi_{\mathrm{min}}-\epsilon_{\mathrm{min}}}\bigg{[}\sqrt{2}+4\lambda(k-1)\frac{\epsilon_{\mathrm{min}}}{\sqrt{k}})\bigg{]}\right\}^{2}}_{\tau}\sum_{j=1}^{k}\|\bm{h}^{t}_{j}\|^{2}_{2}+8k\left[\frac{\tau_{\mathrm{noise}}}{(1-\varrho)\pi_{\min}-\epsilon_{\mathrm{min}}}\right]^{2}$
$=\tau\|\bm{h}^{t}\|^{2}_{2}+C\sigma^{2}_{z}\left(\frac{sk\log\left(\frac{n}{s}\right)+s\log\left(\frac{d}{s}\right)+\log\left(\frac{1}{\delta}\right)}{n}\right),$ (61)
@@
$\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}^{t}_{j}\|_{2}+\|\bm{\widetilde{\Pi}}_{\varphi^{t}\backslash\varphi^{t+1}}\bm{h}^{t}_{j}\|_{2}\leq\sqrt{2}\|\bm{\widetilde{\Pi}}_{\varphi^{t}}\bm{h}^{t}_{j}\|_{2},$ (62)
@@
$\tau^{t+1}\left\|\bm{\theta}^{0}-\bm{\theta}^{\star}\right\|^{2}_{2}\leq(\Delta\rho)^{2}.$ (64)
Next, the second term on the right-hand side of (63) is upper bounded as
@@
$\mathcal{Z}_{s}\triangleq\left\{\varphi\subset[d]:|\varphi|=s\right\}.$
@@
$\mathcal{P}_{k,d,s}=\bigcup_{\varphi\in\mathcal{Z}_{s}}\mathcal{P}_{k,d}\left(\varphi\right),$ (67)
@@
$\mathcal{P}_{k,d}(\varphi)\triangleq\left\{\bm{x}\in\mathbb{R}^{d}:[\bm{x}]_{\varphi^{\mathsf{c}}}=\bm{0},~{}M[\bm{x}]_{\varphi}\geq\bm{b},~{}M\in\mathbb{R}^{k\times s},~{}\bm{b}\in\mathbb{R}^{k}\right\}.$ (68)
We use $\bm{\Pi}_{\varphi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to denote the orthogonal projection onto the subspace spanned by the standard basis vectors $\{\bm{e}_{j}\}_{j\in\varphi}$ for $\varphi\subset[d]$. Then the augmented operator $\bm{\widetilde{\Pi}}_{\varphi}:\mathbb{R}^{d+1}\rightarrow\mathbb{R}^{d+1}$ is defined as
$\bm{\widetilde{\Pi}}_{\varphi}=\left[{\begin{array}[]{cc}\bm{\Pi}_{\varphi}&\bm{0}_{d\times 1}\\ \bm{0}_{1\times d}&1\\ \end{array}}\right].$ (69)
@@
Let $\{\bm{x}_{i}\}_{i=1}^{n}$ be independent copies of a random vector $\bm{x}\in\mathbb{R}^{d}$ which satisfies Assumptions 2.1 and 2.2. Let $\{\omega_{i}\}_{i=1}^{n}\in\{0,1\}^{n}$ be fixed with $\sum_{i=1}^{n}\omega_{i}=\ell>0$. Then for all $\epsilon\in[0,1]$, there exists an absolute constant $C>0$ where it holds with probability at least $1-\delta$ that
$\underset{\varphi\in\mathcal{Z}_{s}}{\mathrm{sup}}{\left\lVert\bm{\widetilde{\Pi}}_{\varphi}\sum_{i=1}^{n}\omega_{i}\left([\bm{x}_{i};1][\bm{x}_{i};1]^{{\scriptscriptstyle{\textup{{T}}}}}-\bm{I}_{d+1}\right)\bm{\widetilde{\Pi}}_{\varphi}\right\rVert}\leq\ell\epsilon$ (70)
@@
The proof of this lemma is obtained by a direct application of the union bound to the statement in [Kim and Lee, 2024, Corollary 6.7]. The result of [Kim and Lee, 2024, Corollary 6.7] provides the concentration inequality in (73) when the supremum is over $\mathcal{P}_{k,d}(\varphi)$ for a fixed $\varphi$ instead of the union over $\varphi\in\mathcal{Z}_{s}$. To apply the union bound argument, we inflate the probability of error by the factor $|\mathcal{Z}_{s}|=$ $\binom{d}{s}$ $\leq\left(\frac{ed}{s}\right)^{s}$. ∎
To bound the closeness of the empirical measure in Lemma B.2 to the expectation with ground truth model parameters, we present the following lemma.
@@
Suppose that $\bm{x}\in\mathbb{R}^{d}$ satisfies Assumptions 2.1 and 2.2. Let $\{\mathcal{C}_{j}(\bm{\theta})\}_{j=1}^{k}$ and $\{\mathcal{C}_{j}(\bm{\theta}^{\star})\}_{j=1}^{k}$ be respectively defined by $\bm{\theta}$ and $\bm{\theta}^{\star}\in\mathbb{R}^{dk}$ according to (12). Let $\mathcal{N}(\bm{\theta}^{\star},\Delta\rho)$ be defined with $R>0$ as in (15) and $\varrho\triangleq CR^{2\zeta}k^{-2(1+\zeta^{-1})}$. If $\bm{\theta}\in\mathcal{N}(\bm{\theta}^{\star},\Delta\rho)$, then we have
@@
Let $\{\bm{x}_{i}\}_{i=1}^{n}$ be independent copies of a random vector $\bm{x}$ which satisfies Assumptions 2.1 and 2.2. Then there exists an absolute constant $C$ for which it holds with probability at least $1-\delta$ that
$\sup\limits_{\begin{subarray}{c}\mathcal{C}\in\mathcal{P}_{k,d,s}\\ \varphi\in\mathcal{Z}_{s}\end{subarray}}\left\|\bm{\widetilde{\Pi}}_{\varphi}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}\}}\left([\bm{x}_{i};1][\bm{x}_{i};1]^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\varphi}\right\|\leq\epsilon$ (76)
@@
Let $\{\bm{x}_{i}\}_{i=1}^{n}$ be independent copies of a random vector $\bm{x}$ satisfying Assumptions 2.1 and 2.2. Let $\{z_{i}\}_{i=1}^{n}$ be i.i.d. sub-Gaussian random variables with zero mean and variance $\sigma^{2}_{z}$, independent of everything else. Then there exists an absolute constant $C$ for which it holds with probability at least $1-\delta$ that
$\sup\limits_{\begin{subarray}{c}\mathcal{C}\in\mathcal{P}_{k,d,s}\\ \varphi\in\mathcal{Z}_{s}\end{subarray}}{\left\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}\}}z_{i}\bm{\widetilde{\Pi}}_{\varphi}[\bm{x}_{i};1]\right\rVert}_{2}\leq C\sigma_{z}\sqrt{\frac{sk\log\left(n/s\right)+s\log\left(d/s\right)+\log\left({1}/{\delta}\right)}{n}}.$ (81)
@@
The proof of lemma follows directly from applying the union bound to the statement in [Kim and Lee, 2024, lemma 8.1, Eq. 46]. Since that statement only considers the supremum over non-sparse polytopes, whereas we consider the supremum over jointly sparse polytopes and sparse vector supports, we must inflate the probability of error $\delta$ by $|\mathcal{Z}_{s}|^{2}=$$\binom{d}{s}^{2}$ $\leq\left(\frac{ed}{s}\right)^{2s}$. ∎
@@
The first lemma aims to show that the empirical moment matrix $\widehat{\bm{M}}$, defined in 2, is not rank deficient. In other words, its first $k$ dominant eigenvectors are a basis for the span of $\{\bm{\theta}_{j}^{\star}\}_{j=1}^{k}$. To state this lemma, we will make use of the definitions in (25) and (24). The first lemma stated next is borrowed from [Ghosh et al., 2021].
@@
$\mathbb{P}(\max_{j\in[k]}|X_{j}|>t)$ $\leq\sum_{j=1}^{k}\mathbb{P}(|X_{j}|>t)\leq\min\left(1,\sum_{j=1}^{k}2\exp\left\{{\frac{-t}{C{\left\lVert X_{j}\right\rVert}}_{\psi_{1}}}\right\}\right),$
$\leq\min\left(1,2k\exp{\frac{-t}{C\underset{j\in[k]}{\max}{\left\lVert X_{j}\right\rVert}_{\psi_{1}}}}\right)=\min\left(1,\exp\left\{\frac{-t}{C\underset{j\in[k]}{\max}{\left\lVert X_{j}\right\rVert}_{\psi_{1}}}+\log 2k\right\}\right),$
$\leq\min\left(1,2\exp{\frac{-t}{C\log_{2}(2k)\underset{j\in[k]}{\max}{\left\lVert X_{j}\right\rVert}_{\psi_{1}}}}\right),$

### v2

@@ 6 Discussion
$\leq 4\sum_{j=1}^{k}\Bigg\{\frac{1}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}\bigg[\epsilon_{\min}\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{h}^{t}_{j}\|_{2}+\epsilon_{\min}\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t}\backslash\mathcal{U}^{t+1}}\bm{h}^{t}_{j}\|_{2}+\lambda\sum_{j^{\prime}=1}^{k}\|\bm{h}^{t}_{j^{\prime}}\|_{2}+\|\bm{d}_{j}\|_{2}\bigg]\Bigg\}^{2}$
$\leq 4\left[\frac{}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}\right]^{2}\sum_{j=1}^{k}\left\{\sqrt{2}\epsilon_{\min}\|\bm{h}^{t}_{j}\|_{2}+\lambda\sum_{j^{\prime}=1}^{k}\|\bm{h}^{t}_{j^{\prime}}\|_{2}+\|\bm{d}_{j}\|_{2}\right\}^{2}$
$\leq\underbrace{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}12}\left(\frac{1}{(1-\varrho)\pi_{\mathrm{min}}-\epsilon_{\mathrm{min}}}\right)^{2}\bigg[2\epsilon^{2}_{\min}+(\lambda k)^{2}\bigg]}_{\tau}\sum_{j=1}^{k}\|\bm{h}^{t}_{j}\|^{2}_{2}+12k\left[\frac{\|\bm{d}_{j}\|_{2}}{(1-\varrho)\pi_{\min}-\epsilon_{\mathrm{min}}}\right]^{2}$
$=\tau\|\bm{h}^{t}\|^{2}_{2}+12k\left[\frac{\|\bm{d}_{1}\|_{2}}{(1-\varrho)\pi_{\min}-\epsilon_{\mathrm{min}}}\right]^{2}$
$=\tau\|\bm{h}^{t}\|^{2}_{2}+\underbrace{C\sigma^{2}_{z}\left(\frac{sk\log\left(\frac{n}{s}\right)+s\log\left(\frac{d}{s}\right)+\log\left(\frac{1}{\delta}\right)}{n}\right)\cdot\frac{k}{\left[(1-\varrho)\pi_{\min}-\epsilon_{\mathrm{min}}\right]^{2}}}_{\rho_{\mathrm{noise}}}$
$=\tau\|\bm{h}^{t}\|^{2}_{2}+\rho_{\mathrm{noise}},$ (70)
@@
$\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{h}^{t}_{j}\|_{2}+\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t}\backslash\mathcal{U}^{t+1}}\bm{h}^{t}_{j}\|_{2}\leq\sqrt{2}\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t}}\bm{h}^{t}_{j}\|_{2},$ (71)
@@
$\tau^{t+1}\left\|\bm{\theta}^{0}-\bm{\theta}^{\star}\right\|^{2}_{2}<\left\|\bm{\theta}^{0}-\bm{\theta}^{\star}\right\|^{2}_{2}\leq(\Delta\rho)^{2}.$ (73)
where the first inequality follows from $\tau<1$ and the second inequality follows from $\bm{\theta}^{0}\in\mathcal{N}(\bm{\theta}^{\star},\Delta\rho)$. Next, the second term on the right-hand side of (72) is upper bounded as
@@
$\mathcal{Z}_{s}\triangleq\left\{\mathcal{U}\subset[d]:|\mathcal{U}|=s\right\}.$
@@
$\mathcal{P}_{k,d,s}=\bigcup_{\mathcal{U}\in\mathcal{Z}_{s}}\mathcal{P}_{k,d}\left(\mathcal{U}\right),$ (76)
@@
$\mathcal{P}_{k,d}(\mathcal{U})\triangleq\left\{\bm{x}\in\mathbb{R}^{d}:[\bm{x}]_{\mathcal{U}^{\mathsf{c}}}=\bm{0},~M[\bm{x}]_{\mathcal{U}}\geq\bm{b},~M\in\mathbb{R}^{k\times s},~\bm{b}\in\mathbb{R}^{k}\right\}.$ (77)
@@
Let $\{\bm{x}_{i}\}_{i=1}^{n}$ be independent copies of a random vector $\bm{x}\in\mathbb{R}^{d}$ which satisfies Assumptions 2.1 and $2.$ Let $\{\omega_{i}\}_{i=1}^{n}\in\{0,1\}^{n}$ be fixed with $\sum_{i=1}^{n}\omega_{i}=\ell>0$. Then for all $\epsilon\in[0,1]$, there exists an absolute constant $C>0$ where it holds with probability at least $1-\delta$ that
$\underset{\mathcal{U}\in\mathcal{Z}_{s}}{\mathrm{sup}}{\left\lVert\bm{\widetilde{\Pi}}_{\mathcal{U}}\sum_{i=1}^{n}\omega_{i}\left([\bm{x}_{i};1][\bm{x}_{i};1]^{{\scriptscriptstyle{\textup{{T}}}}}-\bm{I}_{d+1}\right)\bm{\widetilde{\Pi}}_{\mathcal{U}}\right\rVert}\leq\ell\epsilon$ (78)
@@
The proof of this lemma is obtained by a direct application of the union bound to the statement in [Kim and Lee, 2024, Corollary 6.7]. The result of [Kim and Lee, 2024, Corollary 6.7] provides the concentration inequality in (81) when the supremum is over $\mathcal{P}_{k,d}(\mathcal{U})$ for a fixed $\mathcal{U}$ instead of the union over $\mathcal{U}\in\mathcal{Z}_{s}$. To apply the union bound argument, we inflate the probability of error by the factor $|\mathcal{Z}_{s}|=$ $\binom{d}{s}$ $\leq\left(\frac{ed}{s}\right)^{s}$. ∎
@@
Fix $\delta\in(0,e^{-1})$. Let $\rho$ be defined as in (18) for some $R>0$. Let $\{\mathcal{C}_{j}(\bm{\theta})\}_{j=1}^{k}$ and $\{\mathcal{C}_{j}(\bm{\theta}^{\star})\}_{j=1}^{k}$ be respectively defined by $\bm{\theta}$ and $\bm{\theta}^{\star}\in\mathbb{R}^{dk}$ according to (14). Let $\mathcal{U}\cup\{d+1\}$ denote the joint support of $\{\bm{\theta}_{j}\}_{j=1}^{k}$. Let $\{\bm{x}_{i}\}_{i=1}^{n}$ be independent copies of a random vector $\bm{x}$ satisfying Assumptions 2.1 and $2.$ Then with probability at least $1-\delta$ we have
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}(\bm{\theta})\cap\mathcal{C}_{j^{\prime}}(\bm{\theta}^{\star})\}}\leq\frac{\pi^{2(1+\zeta^{-1})}_{\min}}{k^{2}},$ (83)
if
$n\geq C\left[s\log\left(\frac{d\vee n}{s}\right)+\log\left(\frac{k}{\delta}\right)\right]k^{4}\pi_{\min}^{-4(1+\zeta^{-1})},$ (84)
for all $j^{\prime}\neq j\in[k]$, $\bm{\theta}\in\mathcal{N}(\bm{\theta}^{\star},\Delta\rho)$ and $\mathcal{U}\in\mathcal{Z}_{s}$.
The proof of this lemma is deferred to the Appendix D.3. To bound the closeness of the empirical measure in Lemma B.2 to the expectation with ground truth model parameters, we present the following lemma.
Lemma B.4.
Suppose that $\bm{x}\in\mathbb{R}^{d}$ satisfies Assumptions 2.1 and $2.$ Let $\{\mathcal{C}_{j}(\bm{\theta})\}_{j=1}^{k}$ and $\{\mathcal{C}_{j}(\bm{\theta}^{\star})\}_{j=1}^{k}$ be respectively defined by $\bm{\theta}$ and $\bm{\theta}^{\star}\in\mathbb{R}^{dk}$ according to (14). Let $\mathcal{N}(\bm{\theta}^{\star},\Delta\rho)$ be defined with $R>0$ as in (17) and $\varrho\triangleq CR^{2\zeta}k^{-2(1+\zeta^{-1})}$. If $\bm{\theta}\in\mathcal{N}(\bm{\theta}^{\star},\Delta\rho)$, then we have
@@
Let $\{\bm{x}_{i}\}_{i=1}^{n}$ be independent copies of a random vector $\bm{x}$ which satisfies Assumptions 2.1 and $2.$ Then there exists an absolute constant $C$ for which it holds with probability at least $1-\delta$ that
$\sup\limits_{\begin{subarray}{c}\mathcal{C}\in\mathcal{P}_{k,d,s}\\ \mathcal{U}\in\mathcal{Z}_{s}\end{subarray}}\left\|\bm{\widetilde{\Pi}}_{\mathcal{U}}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}\}}\left([\bm{x}_{i};1][\bm{x}_{i};1]^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\mathcal{U}}\right\|\leq\epsilon$ (86)
@@
Let $\{\bm{x}_{i}\}_{i=1}^{n}$ be independent copies of a random vector $\bm{x}$ satisfying Assumptions 2.1 and $2.$ Let $\{z_{i}\}_{i=1}^{n}$ be i.i.d. sub-Gaussian random variables with zero mean and variance $\sigma^{2}_{z}$, independent of everything else. Then there exists an absolute constant $C$ for which it holds with probability at least $1-\delta$ that
$\sup\limits_{\begin{subarray}{c}\mathcal{C}\in\mathcal{P}_{k,d,s}\\ \mathcal{U}\in\mathcal{Z}_{s}\end{subarray}}{\left\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}\}}z_{i}\bm{\widetilde{\Pi}}_{\mathcal{U}}[\bm{x}_{i};1]\right\rVert}_{2}\leq C\sigma_{z}\sqrt{\frac{sk\log\left(n/s\right)+s\log\left(d/s\right)+\log\left({1}/{\delta}\right)}{n}}.$ (91)
@@
The proof of the lemma follows directly from applying the union bound to the statement in [Kim and Lee, 2024, lemma 8.1, Eq. 46]. Since that statement only considers the supremum over non-sparse polytopes, whereas we consider the supremum over jointly sparse polytopes and sparse vector supports, we must inflate the probability of error $\delta$ by $|\mathcal{Z}_{s}|^{2}=$$\binom{d}{s}^{2}$ $\leq\left(\frac{ed}{s}\right)^{2s}$. ∎
The next lemma provides a tail bound on the worst-case eigenvalue of the sum of covariate outer product with bounded cardinality.
Lemma B.8.
Let $\delta\in(0,e^{-1})$ and $\alpha\in(0,1)$. Let $\{\bm{x}_{i}\}_{i=1}^{n}$ be independent copies of a random vector $\bm{x}$ that satisfies Assumption 2.1. Then, with probability at least $1-\delta$ we have
$\sup_{\begin{subarray}{c}\mathcal{I}:|\mathcal{I}|\leq\alpha n\\ \mathcal{U}\in\mathcal{Z}_{s}\end{subarray}}\lambda_{1}\left[\widetilde{\bm{\Pi}}_{\mathcal{U}}\left(\frac{1}{n}\sum_{i\in\mathcal{I}}[\bm{x}_{i};1][\bm{x}_{i};1]^{\top}\right)\widetilde{\bm{\Pi}}_{\mathcal{U}}^{\top}\right]\leq C(\eta^{2}\vee 1)\sqrt{\alpha},$ (92)
if
$n\geq\alpha^{-1}\left[s\log(d/s)+\log(1/\delta)\right].$ (93)
Proof.
For fixed $\mathcal{U}\in\mathcal{Z}_{s}$, (92) follows directly from [Tan and Vershynin, 2019, Theorem 5.7] if $n\geq d\vee\alpha^{-1}\log(1/\delta)$. Using the union bound and inflating the probability of error by $|\mathcal{Z}_{s}|=$$\binom{d}{s}$ $\leq\left(\frac{ed}{s}\right)^{s}$ completes the proof. ∎
@@
The first lemma aims to show that the empirical moment matrix $\widehat{\bm{M}}$, defined in Algorithm 2, is not rank deficient. In other words, its first $k$ dominant eigenvectors are a basis for the span of $\{\bm{\theta}_{j}^{\star}\}_{j=1}^{k}$. To state this lemma, we will make use of the definitions in (27) and (26). The first lemma stated next is borrowed from [Ghosh et al., 2021].
@@
$\mathbb{P}(\max_{j\in[k]}|X_{j}|>t)$ $\leq\sum_{j=1}^{k}\mathbb{P}(|X_{j}|>t)\leq 1\wedge\sum_{j=1}^{k}2\exp\left\{{\frac{-t}{C{\left\lVert X_{j}\right\rVert}}_{\psi_{1}}}\right\},$
$\leq 1\wedge 2k\exp\left\{\frac{-t}{C\underset{j\in[k]}{\max}{\left\lVert X_{j}\right\rVert}_{\psi_{1}}}\right\}=1\wedge\exp\left\{\frac{-t}{C\underset{j\in[k]}{\max}{\left\lVert X_{j}\right\rVert}_{\psi_{1}}}+\log 2k\right\},$
$\leq 1\wedge 2\exp\left\{\frac{-t}{C\log_{2}(2k)\underset{j\in[k]}{\max}{\left\lVert X_{j}\right\rVert}_{\psi_{1}}}\right\},$